a hose can pump 6 cups of water per second.how many gallons would it pump in one hour? how do i do this math problem?
Answer by dawny74(8)

a hose can pump 6 cups of water per second.how many gallons would it pump in one hour? how do i do this math problem?
First, evaluate -
You will need to determine how many cups are in a gallon - there are 16 cups in a gallon
Next you need to know how many seconds in a hour 60 (seconds in a minute) * 60 (minutes in a hour) = 3600 seconds in a hour.
Now you know that the hose pumps 6 cups per second, but you need to figure out how many gallons it will pump in an hour
a * b = c
Let a = number of cups in a second
Let b = number of seconds in an hour
Let c = total number of cups in a hour
6 * 3600 = c
c= 21,600 cups in an hour
Now you will need to divide c by x (which will equal the number of cups in a gallon)
c/x = z (number of gallons in an hour
21,600/16 = z
z=1350 gallons